OpenStudy (anonymous):

two boys want to use a seesaw, and they need to move the seesaw so that their weight will balance out. the formula is given by w1*d1=w2*d2 where w1= weight of the first boy, d1= distance of the first boy from the fulcrum, w2= weight of the second boy, and d2= distance of the second boy from fulcrum. rewrite the formula to solve for d2. show all steps in your work.

OpenStudy (goformit100):

Welcome to Open Study ........... Sir/Ma'am @Genesisga

OpenStudy (anonymous):

can u please help me with this problem i dont understand it

OpenStudy (goformit100):

w1*d1=w2*d2 or, w2*d2 = w1*d1 or, (w2*d2)/w2 = (w1*d1)w2 or, d2 = (w1*d1)w2 ................................... is the required answer

OpenStudy (anonymous):

thank you @goformit100